x+x+x+1=201
3x+1=201

The sum of the first number is 66, the second 67, and the third number is 68.

Is this correct and please explain !

Answer by checkley71(8403)

X+X+X+1=201
3X=201-1
3X=200
X=66.67 ANSWER.
------------------------------------------------
IF YOU MEANT:
WHAT 3 COSECUTIVE NUMBERS TOTAL 201?
THEN:
X+(X+1)+(X+2)=201
3X+3=201
3X=201-3
3X=198
X=66
X+1=67
X+2=68